# Critical dynamics of Model H within the real-time fRG approach

### v1

@@ Appendix B Loop diagram calculation
In this appendix, we give a specific example to show how the calculation are performed and what approximation are made. We employ a loop diagram for the two-point correlation function $\Gamma_{\phi_{q}\phi_{c},k}^{(2)}$ with the regulator function inserting on the order parameter propagator, that is,
@@
Taking the expression of vertices and propagators in App. A into the above equation, one has
@@
$=$ $\int_{q}\partial_{\tau}R_{\phi,k}g_{k}^{2}(-1+\cos^{2}\theta)\bm{q}^{2}\bm{p}^{2}\frac{4\eta_{k}}{q_{0}^{2}+\left[\eta_{k}\bm{q}^{2}\Big{(}1+r_{B}\Big{(}\frac{\bm{q}^{2}}{k^{2}}\Big{)}\Big{)}\right]^{2}}\frac{1}{\left[iq_{0}+\lambda_{k}\bm{q}^{2}\left(\bm{q}^{2}\Big{(}1+r_{B}\Big{(}\frac{\bm{q}^{2}}{k^{2}}\Big{)}\Big{)}+m_{\phi,k}^{2}\right)\right]^{2}}$ (45)
The vertices keep the external momentum to order 2. Since the flow equation for transport coefficient $\lambda_{k}$ project onto the $p^{4}$ term, we need to expand the $G_{j,k}^{K}$ propagator about the external momentum. The result is given by
$-\frac{1}{2}\frac{1}{[q_{0}^{2}/k^{2}+(\eta_{k}x(1+r_{B}(x)))^{2}]^{2}}2\eta_{k}^{2}\delta(1-x)\frac{4(\bm{q}\cdot\bm{p})^{2}}{k^{4}}$ (46)
@@
Now we complete the integration in Eq. 45,
$\frac{1}{(2\pi)^{d}}\int\mathrm{dx}\mathrm{d\Omega_{D}}k^{d}x^{\frac{d}{2}}16g_{k}^{2}\lambda_{k}\eta_{k}^{3}(-1+\cos^{2}\theta)\cos^{2}\theta\bm{q}^{4}\bm{p}^{4}\frac{1}{k^{16}}\frac{k^{6}(3\eta_{k}+\lambda_{k}k^{2}x(1+\bar{m}^{2}))}{4\eta_{k}^{3}(\eta_{k}+\lambda_{k}k^{2}x(1+\bar{m}^{2}))^{3}}\delta(1-x)$ (47)
Here comes to the most important approximation we made in order to get the final result, the shear viscosity $\eta_{k}$ grows large with the running of RG scale while $\lambda_{k}$ approaches to a finite value in the IR. With the RG scale evolving to IR, we have $\eta_{k}>>\lambda_{k}$. So the $\lambda_{k}k^{2}$ term in the denominator can be neglected and we get the final expression
$\frac{1}{(2\pi)^{d}}\int\mathrm{d\Omega_{D}}\frac{6\bar{g}^{2}}{\bar{\eta}^{2}}(-1+\cos^{2}\theta)\cos^{2}\theta\lambda_{k}\bm{p}^{4}$ (48)

### v2

@@ Appendix B Loop diagram calculation
In this appendix, we present a specific example to show how the calculation are performed and to illustrate the approximations used. We employ a loop diagram for the two-point correlation function $\Gamma_{\phi_{q}\phi_{c},k}^{(2)}$ with the regulator function inserted on the propagator of order parameter, that is,
@@
Substituting the expressions of vertices and propagators in App. A into the equation above, one has
@@
$=$ $\int_{q}\partial_{\tau}R_{\phi,k}g_{k}^{2}(-1+\cos^{2}\theta)\bm{q}^{2}\bm{p}^{2}\frac{4\eta_{k}}{q_{0}^{2}+\left[\eta_{k}\bm{q}^{2}\Big{(}1+r_{B}\Big{(}\frac{\bm{q}^{2}}{k^{2}}\Big{)}\Big{)}\right]^{2}}\frac{1}{\left[iq_{0}+\lambda_{k}\bm{q}^{2}\left(\bm{q}^{2}\Big{(}1+r_{B}\Big{(}\frac{\bm{q}^{2}}{k^{2}}\Big{)}\Big{)}+m_{\phi,k}^{2}\right)\right]^{2}}\,.$ (48)
The vertices contribute to the external momentum of order $p^{2}$. Since the flow equation for transport coefficient $\lambda_{k}$ should be projected onto the $p^{4}$ term, we need to expand the $G_{j,k}^{K}$ propagator about the external momentum. The result is given by
$-\frac{1}{2}\frac{1}{[q_{0}^{2}/k^{2}+(\eta_{k}x(1+r_{B}(x)))^{2}]^{2}}2\eta_{k}^{2}\delta(1-x)\frac{4(\bm{q}\cdot\bm{p})^{2}}{k^{4}}\,,$ (49)
@@
Then we complete the integration in Eq. 48,
$\frac{1}{(2\pi)^{d}}\int\mathrm{d}x\mathrm{d\Omega_{D}}k^{d}x^{\frac{d}{2}}16g_{k}^{2}\lambda_{k}\eta_{k}^{3}(-1+\cos^{2}\theta)\cos^{2}\theta\bm{q}^{4}\bm{p}^{4}\frac{1}{k^{16}}\frac{k^{6}(3\eta_{k}+\lambda_{k}k^{2}x(1+\bar{m}^{2}))}{4\eta_{k}^{3}(\eta_{k}+\lambda_{k}k^{2}x(1+\bar{m}^{2}))^{3}}\delta(1-x)\,.$ (50)
As the RG scale evolves towards the infrared, the shear viscosity $\eta_{k}$ increases, and $\lambda_{k}$ approaches to a finite value. Then we have $\eta_{k}>>\lambda_{k}$, and the $\lambda_{k}k^{2}$ term in the denominator can be safely neglected, which leads us to
$\frac{1}{(2\pi)^{d}}\int\mathrm{d\Omega_{D}}\frac{6\bar{g}^{2}}{\bar{\eta}^{2}}(-1+\cos^{2}\theta)\cos^{2}\theta\lambda_{k}\bm{p}^{4}\,.$ (51)